# Optimal frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v1

@@ 2.1. Notation
For a positive integer $m$, we define $m]:=\{1,\dotsc,m\}$. Let $\mathbb{H}\in\{\mathbb{R},\mathbb{C}\}$. We use $\|\bm{x}\|_{2}$ to denote the Euclidean norm of a vector $\bm{x}\in\mathbb{H}^{d}$, and we use $\|\bm{A}\|_{2}$ to denote the spectral norm of a matrix $\bm{A}\in\mathbb{H}^{m\times d}$. We denote the unit sphere in $\mathbb{H}^{d}$ by $\mathbb{S}_{\mathbb{H}}^{d-1}$, i.e.,
$\mathbb{S}_{\mathbb{H}}^{d-1}=\{\bm{x}\in\mathbb{H}^{d}:\|\bm{x}\|_{2}=1\}.$
@@ 2.2. Equivalent forms of Lipschitz constants and condition numbers
[8, 6, 1] Let $\bm{A}\in\mathbb{H}^{m\times d}$, where $\mathbb{H}=\mathbb{R}$ or $\mathbb{C}$. Then $U_{\bm{A}}=\|\bm{A}\|_{2}$.
@@
[2] Let $\bm{A}\in\mathbb{H}^{m\times d}$, where $\mathbb{H}=\mathbb{R}$ or $\mathbb{C}$. Then
$L_{\bm{A}}=\min_{\begin{subarray}{c}\bm{x},\bm{y}\in\mathbb{H}^{d}\\ \|\bm{x}\|_{2}=1,\|\bm{y}\|_{2}\leq 1,\left<\bm{x},\bm{y}\right>=0\end{subarray}}\frac{\||\bm{A}\bm{x}|-|\bm{A}\bm{y}|\|_{2}}{\operatorname{dist}(\bm{x},\bm{y})}.$
@@
Assume that $\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\subset\mathbb{H}^{d}$ forms a tight frame, where $\mathbb{H}=\mathbb{R}$ or $\mathbb{C}$. Then
$\beta_{\bm{A}}=\left(1-\frac{1}{C}\underset{\begin{subarray}{c}\bm{x},\bm{y}\in\mathbb{H}^{d},\\ \|\bm{x}\|_{2}=\|\bm{y}\|_{2}=1,\langle\bm{x},\bm{y}\rangle=0\end{subarray}}{\max}\sum_{j=1}^{m}\left|\bm{x}^{*}\bm{a}_{j}\bm{a}_{j}^{*}\bm{y}\right|\right)^{-\frac{1}{2}},$ (2.1)
where $\bm{A}=[\bm{a}_{1},\dotsc,\bm{a}_{m}]^{*}\in\mathbb{H}^{m\times d}$ and $C=\frac{1}{d}\sum_{j=1}^{m}{\|\bm{a}_{j}\|_{2}^{2}}$.
@@
Since $\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\subset\mathbb{H}^{d}$ forms a tight frame, we have $\bm{A}^{*}\bm{A}=C\cdot\bm{I}_{d}$, where $C=\frac{1}{d}\sum_{j=1}^{m}{\|\bm{a}_{j}\|_{2}^{2}}$. By Theorem 2.2 we have
$(L_{\bm{A}})^{2}=\min_{\begin{subarray}{c}\bm{x},\bm{y}\in\mathbb{H}^{d}\\ \|\bm{x}\|_{2}=1,\|\bm{y}\|_{2}\leq 1,\left<\bm{x},\bm{y}\right>=0\end{subarray}}\frac{\||\bm{A}\bm{x}|-|\bm{A}\bm{y}|\|_{2}^{2}}{\operatorname{dist}^{2}(\bm{x},\bm{y})}.$ (2.2)
If $\bm{y}=\mathbf{0}$, then we simply have
$\frac{\||\bm{A}\bm{x}|-|\bm{A}\bm{y}|\|_{2}^{2}}{\operatorname{dist}^{2}(\bm{x},\bm{y})}=\frac{\|\bm{A}\bm{x}\|_{2}^{2}}{\|\bm{x}\|_{2}^{2}}=C.$ (2.3)
We next consider the case when $\bm{y}\neq\mathbf{0}$. Assume that $\bm{x},\,\bm{y}\in\mathbb{H}^{d}$, $\|\bm{x}\|_{2}=1,\,0<\|\bm{y}\|_{2}\leq 1$, and $\left<\bm{x},\bm{y}\right>=0$. A direct calculation shows $\operatorname{dist}^{2}(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|_{2}^{2}=1+\|\bm{y}\|_{2}^{2}$ and
@@
where $\widetilde{\bm{y}}:=\bm{y}/\|\bm{y}\|_{2}$. It follows that
@@
Here, we apply the Cauchy-Schwarz inequality in ($a$), where the equality is achieved if and only if $\|\bm{y}\|_{2}=1$. Therefore, combining (2.2) with (2.3) and (2.4), we have
$(L_{\bm{A}})^{2}=\underset{\begin{subarray}{c}\bm{x},\bm{y}\in\mathbb{H}^{d},\\ \|\bm{x}\|_{2}=\|\bm{y}\|_{2}=1,\,\langle\bm{x},\bm{y}\rangle=0\end{subarray}}{\min}\left(C-\sum_{j=1}^{m}\left|\bm{x}^{*}\bm{a}_{j}\bm{a}_{j}^{*}\bm{y}\right|\right)=C-\underset{\begin{subarray}{c}\bm{x},\bm{y}\in\mathbb{H}^{d},\\ \|\bm{x}\|_{2}=\|\bm{y}\|_{2}=1,\,\langle\bm{x},\bm{y}\rangle=0\end{subarray}}{\max}\sum_{j=1}^{m}\left|\bm{x}^{*}\bm{a}_{j}\bm{a}_{j}^{*}\bm{y}\right|.$ (2.5)
Combining (2.5) with $\beta_{\bm{A}}=U_{\bm{A}}/L_{\bm{A}}$ and $U_{\bm{A}}=\|\bm{A}\|_{2}=\sqrt{C}$, we arrive at (2.1). ∎
The following lemma establishes that, for determining the minimal condition number of $\bm{A}\in\mathbb{H}^{m\times d}$, it suffices to consider cases where its rows form a tight frame.
@@
$\min_{\bm{A}\in\mathbb{H}^{m\times d}}\beta_{\bm{A}}=\min_{\bm{A}\in\mathbb{H}^{m\times d},\bm{A}^{*}\bm{A}=\bm{I}_{d}}\beta_{\bm{A}}.$
@@
If ${\rm rank}({\bm{A}})<d$, then $\beta_{\bm{A}}=+\infty$. We next consider the case when ${\bm{A}}\in\mathbb{H}^{m\times d}$ has rank $d$. Let $\bm{M}=\bm{A}^{*}\bm{A}\in\mathbb{H}^{d\times d}$ and $\bm{B}=\bm{A}\bm{M}^{-\frac{1}{2}}\in\mathbb{H}^{m\times d}$. We claim that
$\beta_{{\bm{A}}}\geq\beta_{{\bm{B}}}.$ (2.6)
Note that the rows of $\bm{B}$} form a tight frame. Also note that the condition number $\beta_{\bm{B}}$ is invariant under the scaling of $\bm{B}$}. Thus, (2.6) implies the conclusion. It suffices to prove (2.6).
Note that for any $\bm{x},\bm{y}\in\mathbb{H}^{d}$,
$\||\bm{B}\bm{x}|-|\bm{B}\bm{y}|\|_{2}$ $=\||\bm{A}\bm{M}^{-\frac{1}{2}}\bm{x}|-|\bm{A}\bm{M}^{-\frac{1}{2}}\bm{y}|\|_{2}=\||\bm{A}\widehat{\bm{x}}|-|\bm{A}\widehat{\bm{y}}|\|_{2},$
where $\widehat{\bm{x}}=\bm{M}^{-\frac{1}{2}}\bm{x}$ and $\widehat{\bm{y}}=\bm{M}^{-\frac{1}{2}}\bm{y}$. Also note that for any $c\in\mathbb{H}$,
$\|\bm{x}-c\cdot\bm{y}\|_{2}=\|\bm{M}^{\frac{1}{2}}(\widehat{\bm{x}}-c\cdot\widehat{\bm{y}})\|_{2}\leq\|\bm{M}^{\frac{1}{2}}\|_{2}\cdot\|\widehat{\bm{x}}-c\cdot\widehat{\bm{y}}\|_{2},$
@@
$\frac{\||\bm{B}\bm{x}|-|\bm{B}\bm{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})}=\frac{\||\bm{A}\widehat{\bm{x}}|-|\bm{A}\widehat{\bm{y}}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})}\geq\frac{1}{\|\bm{M}^{\frac{1}{2}}\|_{2}}\cdot\frac{\||\bm{A}\widehat{\bm{x}}|-|\bm{A}\widehat{\bm{y}}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\widehat{\bm{x}},\widehat{\bm{y}})}$.
Since $\bm{M}$ has full rank, we have
$L_{\bm{B}}=\inf_{\begin{subarray}{c}\bm{x},\bm{y}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})\neq 0\end{subarray}}\frac{\||\bm{B}\bm{x}|-|\bm{B}\bm{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})}\geq\frac{1}{\|\bm{M}^{\frac{1}{2}}\|_{2}}\cdot\inf_{\begin{subarray}{c}\widehat{\bm{x}},\widehat{\bm{y}}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\widehat{\bm{x}},\widehat{\bm{y}})\neq 0\end{subarray}}\frac{\||\bm{A}\widehat{\bm{x}}|-|\bm{A}\widehat{\bm{y}}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\widehat{\bm{x}},\widehat{\bm{y}})}=\frac{L_{\bm{A}}}{\|\bm{M}^{\frac{1}{2}}\|_{2}}.$
Note that ${\bm{B}}^{*}{\bm{B}}=\bm{M}^{-\frac{1}{2}}\bm{A}^{*}\bm{A}\bm{M}^{-\frac{1}{2}}=\bm{I}_{d}$ and $\|\bm{M}^{\frac{1}{2}}\|_{2}=\|\bm{A}\|_{2}$, so we have $U_{\bm{B}}=\|\bm{B}\|_{2}=1$ and
$\beta_{{\bm{B}}}=\frac{U_{\bm{B}}}{L_{\bm{B}}}\leq\frac{\|\bm{M}^{\frac{1}{2}}\|_{2}}{L_{\bm{A}}}=\frac{\|\bm{A}\|_{2}}{L_{\bm{A}}}=\frac{U_{\bm{A}}}{L_{\bm{A}}}=\beta_{{\bm{A}}}.$

### v2

@@ 2.1. Notation
For a positive integer $m$, we define $m]:=\{1,\dotsc,m\}$. Let $\mathbb{H}\in\{\mathbb{R},\mathbb{C}\}$. We use $\|\boldsymbol{x}\|_{2}$ to denote the Euclidean norm of a vector $\boldsymbol{x}\in\mathbb{H}^{d}$, and we use $\|\boldsymbol{A}\|_{2}$ to denote the spectral norm of a matrix $\boldsymbol{A}\in\mathbb{H}^{m\times d}$. We denote the unit sphere in $\mathbb{H}^{d}$ by $\mathbb{S}_{\mathbb{H}}^{d-1}$, i.e.,
$\mathbb{S}_{\mathbb{H}}^{d-1}=\{\boldsymbol{x}\in\mathbb{H}^{d}:\|\boldsymbol{x}\|_{2}=1\}.$
@@ 2.2. Equivalent forms of Lipschitz constants and condition numbers
[8, 6, 1] Let $\boldsymbol{A}\in\mathbb{H}^{m\times d}$, where $\mathbb{H}=\mathbb{R}$ or $\mathbb{C}$. Then $U_{\boldsymbol{A}}=\|\boldsymbol{A}\|_{2}$.
@@
[2] Let $\boldsymbol{A}\in\mathbb{H}^{m\times d}$, where $\mathbb{H}=\mathbb{R}$ or $\mathbb{C}$. Then
$L_{\boldsymbol{A}}=\min_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}\\ \|\boldsymbol{x}\|_{2}=1,\|\boldsymbol{y}\|_{2}\leq 1,\left<\boldsymbol{x},\boldsymbol{y}\right>=0\end{subarray}}\frac{\||\boldsymbol{A}\boldsymbol{x}|-|\boldsymbol{A}\boldsymbol{y}|\|_{2}}{\operatorname{dist}(\boldsymbol{x},\boldsymbol{y})}.$
@@
Assume that $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{H}^{d}$ forms a tight frame, where $\mathbb{H}=\mathbb{R}$ or $\mathbb{C}$. Then
$\beta_{\boldsymbol{A}}=\left(1-\frac{1}{C}\underset{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d},\\ \|\boldsymbol{x}\|_{2}=\|\boldsymbol{y}\|_{2}=1,\langle\boldsymbol{x},\boldsymbol{y}\rangle=0\end{subarray}}{\max}\sum_{j=1}^{m}\left|\boldsymbol{x}^{*}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{*}\boldsymbol{y}\right|\right)^{-\frac{1}{2}},$ (2.1)
where $\boldsymbol{A}=[\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}]^{*}\in\mathbb{H}^{m\times d}$ and $C=\frac{1}{d}\sum_{j=1}^{m}{\|\boldsymbol{a}_{j}\|_{2}^{2}}$.
@@
Since $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{H}^{d}$ forms a tight frame, we have $\boldsymbol{A}^{*}\boldsymbol{A}=C\cdot\boldsymbol{I}_{d}$, where $C=\frac{1}{d}\sum_{j=1}^{m}{\|\boldsymbol{a}_{j}\|_{2}^{2}}$. By Theorem 2.2 we have
$(L_{\boldsymbol{A}})^{2}=\min_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}\\ \|\boldsymbol{x}\|_{2}=1,\|\boldsymbol{y}\|_{2}\leq 1,\left<\boldsymbol{x},\boldsymbol{y}\right>=0\end{subarray}}\frac{\||\boldsymbol{A}\boldsymbol{x}|-|\boldsymbol{A}\boldsymbol{y}|\|_{2}^{2}}{\operatorname{dist}^{2}(\boldsymbol{x},\boldsymbol{y})}.$ (2.2)
If $\boldsymbol{y}=\mathbf{0}$, then we simply have
$\frac{\||\boldsymbol{A}\boldsymbol{x}|-|\boldsymbol{A}\boldsymbol{y}|\|_{2}^{2}}{\operatorname{dist}^{2}(\boldsymbol{x},\boldsymbol{y})}=\frac{\|\boldsymbol{A}\boldsymbol{x}\|_{2}^{2}}{\|\boldsymbol{x}\|_{2}^{2}}=C.$ (2.3)
We next consider the case when $\boldsymbol{y}\neq\mathbf{0}$. Assume that $\boldsymbol{x},\,\boldsymbol{y}\in\mathbb{H}^{d}$, $\|\boldsymbol{x}\|_{2}=1,\,0<\|\boldsymbol{y}\|_{2}\leq 1$, and $\left<\boldsymbol{x},\boldsymbol{y}\right>=0$. A direct calculation shows $\operatorname{dist}^{2}(\boldsymbol{x},\boldsymbol{y})=\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2}=1+\|\boldsymbol{y}\|_{2}^{2}$ and
$\left\||\boldsymbol{A}\boldsymbol{x}|-|\boldsymbol{A}\boldsymbol{y}|\right\|_{2}^{2}$ $=\boldsymbol{x}^{*}\boldsymbol{A}^{*}\boldsymbol{A}\boldsymbol{x}+\boldsymbol{y}^{*}\boldsymbol{A}^{*}\boldsymbol{A}\boldsymbol{y}-2\sum_{j=1}^{m}\left|\boldsymbol{x}^{*}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{*}\boldsymbol{y}\right|$
@@
where $\widetilde{\boldsymbol{y}}:=\boldsymbol{y}/\|\boldsymbol{y}\|_{2}$. It follows that
@@
Here, we apply the Cauchy-Schwarz inequality in ($a$), where the equality is achieved if and only if $\|\boldsymbol{y}\|_{2}=1$. Therefore, combining (2.2) with (2.3) and (2.4), we have
$(L_{\boldsymbol{A}})^{2}$ $=\underset{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d},\\ \|\boldsymbol{x}\|_{2}=\|\boldsymbol{y}\|_{2}=1,\,\langle\boldsymbol{x},\boldsymbol{y}\rangle=0\end{subarray}}{\min}\left(C-\sum_{j=1}^{m}\left|\boldsymbol{x}^{*}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{*}\boldsymbol{y}\right|\right)$ (2.5)
$=C-\underset{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d},\\ \|\boldsymbol{x}\|_{2}=\|\boldsymbol{y}\|_{2}=1,\,\langle\boldsymbol{x},\boldsymbol{y}\rangle=0\end{subarray}}{\max}\sum_{j=1}^{m}\left|\boldsymbol{x}^{*}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{*}\boldsymbol{y}\right|.$
Combining (2.5) with $\beta_{\boldsymbol{A}}=U_{\boldsymbol{A}}/L_{\boldsymbol{A}}$ and $U_{\boldsymbol{A}}=\|\boldsymbol{A}\|_{2}=\sqrt{C}$, we arrive at (2.1). ∎
The following lemma establishes that, for determining the minimal condition number of $\boldsymbol{A}\in\mathbb{H}^{m\times d}$, it suffices to consider cases where its rows form a tight frame.
@@
$\min_{\boldsymbol{A}\in\mathbb{H}^{m\times d}}\beta_{\boldsymbol{A}}=\min_{\boldsymbol{A}\in\mathbb{H}^{m\times d},\boldsymbol{A}^{*}\boldsymbol{A}=\boldsymbol{I}_{d}}\beta_{\boldsymbol{A}}.$
@@
If ${\rm rank}({\boldsymbol{A}})<d$, then $\beta_{\boldsymbol{A}}=+\infty$. We next consider the case when ${\boldsymbol{A}}\in\mathbb{H}^{m\times d}$ has rank $d$. Let $\boldsymbol{M}=\boldsymbol{A}^{*}\boldsymbol{A}\in\mathbb{H}^{d\times d}$ and $\boldsymbol{B}=\boldsymbol{A}\boldsymbol{M}^{-\frac{1}{2}}\in\mathbb{H}^{m\times d}$. We claim that
$\beta_{{\boldsymbol{A}}}\geq\beta_{{\boldsymbol{B}}}.$ (2.6)
Note that the rows of $\boldsymbol{B}$} form a tight frame. Also note that the condition number $\beta_{\boldsymbol{B}}$ is invariant under the scaling of $\boldsymbol{B}$}. Thus, (2.6) implies the conclusion. It suffices to prove (2.6).
Note that for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}$,
$\||\boldsymbol{B}\boldsymbol{x}|-|\boldsymbol{B}\boldsymbol{y}|\|_{2}$ $=\||\boldsymbol{A}\boldsymbol{M}^{-\frac{1}{2}}\boldsymbol{x}|-|\boldsymbol{A}\boldsymbol{M}^{-\frac{1}{2}}\boldsymbol{y}|\|_{2}=\||\boldsymbol{A}\widehat{\boldsymbol{x}}|-|\boldsymbol{A}\widehat{\boldsymbol{y}}|\|_{2},$
where $\widehat{\boldsymbol{x}}=\boldsymbol{M}^{-\frac{1}{2}}\boldsymbol{x}$ and $\widehat{\boldsymbol{y}}=\boldsymbol{M}^{-\frac{1}{2}}\boldsymbol{y}$. Also note that for any $c\in\mathbb{H}$,
$\|\boldsymbol{x}-c\cdot\boldsymbol{y}\|_{2}=\|\boldsymbol{M}^{\frac{1}{2}}(\widehat{\boldsymbol{x}}-c\cdot\widehat{\boldsymbol{y}})\|_{2}\leq\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}\cdot\|\widehat{\boldsymbol{x}}-c\cdot\widehat{\boldsymbol{y}}\|_{2},$
@@
$\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})$ $=\min_{c\in\mathbb{H},|c|=1}\|\boldsymbol{x}-c\cdot\boldsymbol{y}\|_{2}$
@@
$\frac{\||\boldsymbol{B}\boldsymbol{x}|-|\boldsymbol{B}\boldsymbol{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})}=\frac{\||\boldsymbol{A}\widehat{\boldsymbol{x}}|-|\boldsymbol{A}\widehat{\boldsymbol{y}}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})}\geq\frac{1}{\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}}\cdot\frac{\||\boldsymbol{A}\widehat{\boldsymbol{x}}|-|\boldsymbol{A}\widehat{\boldsymbol{y}}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}})}$.
Since $\boldsymbol{M}$ has full rank, we have
$L_{\boldsymbol{B}}$ $=\inf_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})\neq 0\end{subarray}}\frac{\||\boldsymbol{B}\boldsymbol{x}|-|\boldsymbol{B}\boldsymbol{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})}$
$\geq\frac{1}{\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}}\cdot\inf_{\begin{subarray}{c}\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}})\neq 0\end{subarray}}\frac{\||\boldsymbol{A}\widehat{\boldsymbol{x}}|-|\boldsymbol{A}\widehat{\boldsymbol{y}}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}})}=\frac{L_{\boldsymbol{A}}}{\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}}.$
Note that ${\boldsymbol{B}}^{*}{\boldsymbol{B}}=\boldsymbol{M}^{-\frac{1}{2}}\boldsymbol{A}^{*}\boldsymbol{A}\boldsymbol{M}^{-\frac{1}{2}}=\boldsymbol{I}_{d}$ and $\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}=\|\boldsymbol{A}\|_{2}$, so we have $U_{\boldsymbol{B}}=\|\boldsymbol{B}\|_{2}=1$ and
$\beta_{{\boldsymbol{B}}}=\frac{U_{\boldsymbol{B}}}{L_{\boldsymbol{B}}}\leq\frac{\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}}{L_{\boldsymbol{A}}}=\frac{\|\boldsymbol{A}\|_{2}}{L_{\boldsymbol{A}}}=\frac{U_{\boldsymbol{A}}}{L_{\boldsymbol{A}}}=\beta_{{\boldsymbol{A}}}.$